XYZ corporation has two employees that cost a total of $90,000 annually. The two employees are given $26,000,000 in A/R. On 12/31/18 the two employees collected $3,110,000 and wrote of $22,890,000. Better collectors inc., an outsourced vendor states they can take the same $26,000,000 in A/R and collect 5,241,600 and only write of 20,758,400. The fee to Better collectors inc would be 13% of what they collected (5,241,600 * 13% = 681,408).

Should XYZ corporation terminate or repurpose the two employees and give the business to Better collectors inc.? What ROI would Better collectors inc state to XYZ corporation.

Please give detailed response.
Answer #1

If 2 employees are used: Scenario 1:

Cost = $90000

Benefit = collection of 3110000$

Net benefit = 3110000-90000=3020000$

Scenario 2:

Cost = fee to better collectors inc = 5,241,600 * 13% = 681,408

Benefit =5,241,600 $collections

Net benefit = 5,241,600-681,408=4560192$

Therefore its better to go with better collectors inc.

Incremental benefit of going ahead with scenario 2 = 4560192$-3020000$=$1540192

Cost in scenario 2 = 681,408

ROI of scenario 2 = incremental benefit / cost = $1540192/681,408=226%